consider the following information : consumption C=800+0.8Yd, tax rate t=30%, investment I=600, Government Expenditure G=800, Export X= 400, Import I=200+0.2Y, where Y= national Income and Yd= disposable Income

Find the equilibrium Income

Find the value of multiplier

Find the value of consumption at Equilibrium

What is the value of import at equilibrium

What is the value of savings at equilibrium

What is the value of budget surplus

What is the value of current account

Other things remaining the same if investment increases from 600 to 664, the change in equilibrium income is

Other things remaining the same , if tax rate increases to 50%, then the new multiplier is




1
Expert's answer
2021-12-05T18:57:23-0500

Solution:

Equilibrium income: Y = AD

Y = C + I + G + X – M

C = 800 +0.8Yd = 800 + 0.8(Y – T) = 800 + 0.8(Y – 0.3Y)

Y = 800 + 0.8(Y – 0.3Y) + 600 + 800 + (400 – (200+0.2Y))

Y = 800 + 0.8Y – 0.24Y + 600 + 800 + 400 – 200 + 0.2Y

Y = 800 + 600 + 800 + 400 – 200 + 0.8Y – 0.24Y – 0.2Y

Y = 2,400 + 0.36Y

Y – 0.36Y = 2,400

0.64Y = 2,400

Y = 3,750

Equilibrium income = 3,750

 

Multiplier = "\frac{1}{(1 - MPC} = \frac{1}{(1 - 0.8)} = \frac{1}{0.2} = 5"

Multiplier = 5

 

Consumption at equilibrium:

C = 800 + 0.8(Y – 0.3Y) = 800 + 0.8(3,750 – 0.3(3,750)) = 800 + 0.8(3,750 – 1,125)

= 800 + 0.8(2,625) = 800 + 2,100 = 2,900

Consumption at equilibrium = 2,900

 

Value of import at equilibrium:

Import = 200+0.2Y = 200 + 0.2(3,750) = 200 + 750 = 950

Value of import at equilibrium = 950

 

Value of savings at equilibrium:

S = Y – C = 3,750 – 2,900 = 850

Value of savings at equilibrium = 850

 

Value of budget surplus = T – G =

T = 0.3(3,750) = 1,125

G = 800

Budget surplus = 1,125 – 800 = 325

Value of budget surplus = 325


Value of current account:

Current Account = (X – M) + NY + NCT

= (400 – 950) + 0 + 0 = -550

Value of current account = (550)

 

New equilibrium income:

Y = C + I + G + X – M

C = 800 +0.8Yd = 800 + 0.8(Y – T) = 800 + 0.8(Y – 0.3Y)

Y = 800 + 0.8(Y – 0.3Y) + 664 + 800 + (400 – (200+0.2Y))

Y = 800 + 0.8Y – 0.24Y + 664 + 800 + 400 – 200 + 0.2Y

Y = 800 + 664 + 800 + 400 – 200 + 0.8Y – 0.24Y – 0.2Y

Y = 2,464 + 0.36Y

Y – 0.36Y = 2,464

0.64Y = 2,464

Y = 3,850

New equilibrium income = 3,850

The change in equilibrium income = 3,850 – 3,750 = 100


Multiplier = "\frac{-MPC}{(1 - MPC} = \frac{-0.8}{(1 - 0.8)} = \frac{-0.8}{0.2} = -4"

Multiplier = -4